I can't get Cramer's Rule to work for Y. It's OK for X and Z.
the matrix is:
1X + 3Y -1Z = 1
-2X -6Y + 1Z = -3
3X + 5Y -2Z = 4
The Determinate is 2.
X = 2, Y = 1, Z = 1 using Cramer's Rule, but this doesn't check
Using RREF, the solution is 2,0,1 which works.
When I take the determinate of Y in Cramer's rule, it is not 0 which it has to be to make Y=0.
what am I doing wrong?
could it be that Cramer's Rule doesn't work some times?
Thanks
Photonjohn

Answer by stanbon(48569)

I can't get Cramer's Rule to work for Y. It's OK for X and Z.
the matrix is:
1X + 3Y -1Z = 1
-2X -6Y + 1Z = -3
3X + 5Y -2Z = 4
The Determinate is 2.
X = 2, Y = 1, Z = 1 using Cramer's Rule, but this doesn't check
Using RREF, the solution is 2,0,1 which works.
When I take the determinate of Y in Cramer's rule, it is not 0 which it has to be to make Y=0.
what am I doing wrong?
could it be that Cramer's Rule doesn't work some times?
Thanks
----------------------
Cramer's Rule always works.
If the coefficient determinant is zero it means there is no
x,y,z point that satisfies the linear systerm.
----------------------
You have to be careful when using Cramer as a single arithmetic
error will result in a wrong answer.
===========================================
By the way, the answer is x = 2 ; y = 0, z = 1
AND the Y-determinant is zero. Check your
arithmetic.
=======================================
Cheers,
Stan H.